A​ penny, a​ nickel, a​ dime, and a quarter are tossed. What is the probability of obtaining four tails on the​ tosses?
Found 2 solutions by ikleyn, jim_thompson5910:
Answer by ikleyn(52921)
Answer by jim_thompson5910(35256)


As a fraction, that probability is 1/16 because:
(1/2)*(1/2)*(1/2)*(1/2) = 1/(2*2*2*2) = 1/16
The four copies of "1/2" correspond to the probabilities of landing on tails for each individual coin. The multiplication of these "1/2" terms is only possible due to the fact each coin is indepedent of the other three.

If you want to convert to decimal form, then use your calculator to find that 1/16 = 0.0625 which converts to 6.25%